Emily is 41 years old. Colin is 10 years older than Emily. Dan is 15 years younger than Emily. What is the total of their combined ages?

Found 2 solutions by josgarithmetic, saw:
Answer by josgarithmetic(39620)

PERSON     AGE
Emily      41
Colin      41+10
Dan        41-15
TOTAL      very simple arithmetic

Answer by saw(34)

Emily = 41
Colin = 41+10 = 51
Dan = 41-15 = 26
Total of their combined ages = 118